What is the general term of 4,10,18,28,40?

Found 2 solutions by rothauserc, richwmiller:
Answer by rothauserc(4718)

This is a quadratic sequence of order 2
4 10 18 28 40
6 8 10 12
2 2 2
for nth term we know we have a n^2 term
4 10 18 28 40
nth 1 2 3 4 5
n^2 1 4 9 16 25
3 6 9 12 15
therefore the nth term is
nth = n^2 +3n

Answer by richwmiller(17219)

1,1,1,1,1,4
16,8,4,2,1,10
81,27,9,3,1,18
256,64,16,4,1,28
625,125,25,5,1,40
1,0,0,0,0,0
0,1,0,0,0,0
0,0,1,0,0,1
0,0,0,1,0,3
0,0,0,0,1,0
T=a*n^4+b*n^3+c*n^2+dn+e
T=0*n^4+0*n^3+1*n^2+3n+0
T=n^2+3n